OpenStudy (anonymous):

(-2x^+3x)-(5x^-9x+2) What is the first step to solve this equation ?

OpenStudy (anonymous):

add like terms

OpenStudy (anonymous):

Thank You !

OpenStudy (anonymous):

pls be trustworthy with ur concepts!! Example The factors of x2 + 3x - 4 are: (x+4) and (x-1) Why? Well, let us multiply them to see: (x+4)(x-1) = x(x-1) + 4(x-1) = x2 - x + 4x - 4 = x2 + 3x - 4 Multiplying (x+4)(x-1) together is called Expanding. In fact, Expanding and Factoring are opposites: